can i please have help solving a word promblem i have for algebra. the promblem is worded like so: The sum of two numbers is 53. The first number is 137 more than the second number.Find the numbers.
I have found numbers that equal to 53 but i can't get the first number to be 137 more than the second.

Answer by scott8148(6628)

let x=1st and y=2nd

"The sum of two numbers is 53" __ x+y=53

"The first number is 137 more than the second number" __ x=y+137

substituting __ (y+137)+y=53 __ 2y+137=53 __2y=-84 __ y=-42

substituting __ x+(-42)=53 __ x=95